### v1

@@ Abstract
Virtual Reality (VR) and Augmented Reality (AR) technologies are increasingly being integrated into cognitively demanding work environments, such as aviation, surgery, and complex manufacturing processes. These Extended Reality (XR) technologies offer immersive training opportunities, real-time data visualization, and enhanced decision-making support. However, despite its advantages, prolonged exposure to XR can introduce physiological and cognitive challenges, including motion sickness, fatigue, and stress responses, which may impact performance and well-being [52, 57]. Understanding how these environments affect users is crucial for optimizing their application in high-stakes professional settings.
@@
A key issue in immersive environments is cybersickness, a form of motion sickness induced by sensory discrepancies between visual and vestibular signals [53]. It manifests as nausea, dizziness, headache, and other autonomic symptoms frequently experienced by VR users [10, 31]. Motion sickness in VR can occur under various conditions, including prolonged exposure, high visual flow, latency between head movements and visual updates, and discrepancies between acceleration cues and expected motion [68]. The severity and duration of symptoms vary among individuals, with some experiencing mild discomfort that subsides within minutes, while others report persistent symptoms lasting hours after VR exposure [39]. Factors such as individual susceptibility, prior VR experience, and display technology play significant roles in determining the intensity of motion sickness effects.
@@
The washout phase of sickness after XR exposure has been studied, though not in great depth. However, it remains a critical aspect to explore further, particularly given its implications for both the robustness of scientific methods [85] and safety in high-stakes environments. For instance, understanding how long symptoms persist and how quickly they subside is essential when considering the use of XR technologies in professions where precise performance is required, such as surgery or aviation. Prolonged exposure to immersive technologies may lead to lingering symptoms that could impair a user’s ability to operate potentially dangerous machinery or perform high-risk tasks. As XR technologies become more prevalent in work environments, studying recovery from sickness is vital to ensure safe and effective usage in critical settings.
@@
Previous studies have demonstrated that cybersickness can elicit a stress response, reflected in autonomic and endocrine changes. Research has shown that individuals experiencing higher discomfort in VR environments exhibit increased cortisol levels, altered Electrodermal Activity (EDA), and Heart Rate (HR) patterns [6, 50]. A recent study by [34] further supports this by demonstrating that cybersickness caused by Head-Mounted Display (HMD)-based VR environments significantly influences physiological responses, including elevated Heart Rate Variability (HRV), altered autonomic nervous system activity, and increased cortisol levels. The physiological stress response is characterized by the activation of the Hypothalamic-Pituitary-Adrenal (HPA) axis and the sympathetic nervous system [46]. While the HPA axis regulates cortisol release, the sympathetic nervous system governs rapid autonomic adaptations, manifesting in increased EDA and changes in HR [76].
@@
Stress has been shown to impact cognitive functions, particularly working memory. While acute stress can enhance cognitive performance under specific conditions [58], extensive research suggests that high stress levels impair working memory performance [60, 64]. This impairment may be due to increased cortisol and norepinephrine secretion, which modulate neural networks in the prefrontal cortex, thereby affecting working memory function [4].
@@
This study extends previous research by incorporating multiple physiological markers, allowing for a more comprehensive understanding of the motion sickness stress response in VR environments. By examining these effects in a controlled experimental setting, this research contributes to a deeper understanding of the interaction between VR, motion sickness, stress, and cognitive performance.
@@ 2.1 Temporal Progression of Cybersickness During Exposure
Several studies indicate that cybersickness tends to increase progressively with prolonged exposure. [29] observed a linear rise in symptoms among fighter pilots exposed to VR conditions over three hours. Similarly, [63] reported an upward trend in symptoms over a 16-minute duration, while [56] found a comparable linear increase within the first ten minutes. [47] noted a similar trajectory in symptoms over approximately six minutes. [84], who conducted self-reported assessments across two 15-minute conditions, also found indications of a continuous increase.
Longer-duration studies reinforce this trend. [69] examined exposure durations of 15, 30, 45, and 60 minutes in a between-subjects design, identifying a sustained increase in cybersickness beyond the 15-minute mark. [38] further corroborated this finding, showing that longer exposures correlated with more intense symptoms.
A meta-analysis by [59] categorized studies into exposure groups of 0-10 minutes, 10-20 minutes, and beyond 20 minutes. Their findings indicated stable oculomotor symptoms across these groups, whereas nausea and disorientation peaked between 10 and 20 minutes. However, this contrasts with other research suggesting a sustained increase in symptoms beyond 20 minutes.
Some studies suggest adaptation effects over prolonged exposure. In a two-hour experiment, symptoms plateaued after 75 minutes, hinting at potential habituation [9]. [48] found no significant Simulator Sickness Questionnaire [28] (SSQ) score differences between early and later exposure stages, reinforcing the notion of early adaptation. Similarly, studies using the Fast Motion Sickness Questionnaire [30] (FMS) values [47, 74] primarily within the first ten minutes reported adaptation trends.
[67] investigated cybersickness in treadmill-based virtual experiences, noting a significant increase in SSQ disorientation scores up to 15 minutes, followed by an adaptation phase at 45 minutes. Interestingly, nausea peaked at 45 minutes, aligning with [21]’s findings, whereas oculomotor symptoms plateaued after 15 minutes. Their data suggest a more complex progression where adaptation occurs alongside symptom escalation.
Other research challenges the notion of a purely gradual increase in symptoms, instead proposing an episodic pattern [79, 81, 73]. Such fluctuations may be difficult to detect due to limitations in sampling rates, which are often constrained by practical feasibility.
@@ 2.2 Temporal Progression of Cybersickness Post Stimulation
Research is sparse on how cybersickness reacts post-stimulation and shows significant interindividual differences in this aspect. For instance, while most subjects experience minimal symptoms after a 10-60 minute break, some continue to suffer from severe symptoms [66].
A key issue in cybersickness research is the lack of uniformity in post-exposure studies. The need for standard methodologies in this area has been discussed, with suggestions for establishing specific standards [14].
In a study measuring SSQ 30 minutes after VR exposure, no significant differences were found compared to a pre-study baseline, suggesting rapid recovery in most subjects. However, the disorientation subscale remained significantly higher than pre-experiment levels [66]. Contrastingly, [17]’s study indicates more prolonged effects of cybersickness. They measured cybersickness before, immediately after, and one, two, and three hours post-exposure over three days. SSQ values were significantly higher post-exposure each day, challenging the notion of rapid recovery suggested by Singer et al. Yet, both studies found significantly lower SSQ values in the first follow-up compared to immediately post-exposure. Further, recovery patterns vary with the VR experiment’s format. In experiments with high SSQ scores, some participants needed over 30 minutes to recover. In contrast, others reported no discomfort just five minutes post-experiment [72]. This observation was mirrored by [71], where most of the participants no longer felt any symptoms 40 minutes after exposure, whereas 14% still felt strong after-effects.
Few studies have examined participants’ well-being the day after the experiment. For example, a study with pilots post-simulator training revealed that 4.6% reported symptoms 24 hours after training [77]. Another study with pilots showed that those with intense symptoms had significantly higher values two hours post-experiment compared to a less affected group, but no differences were observed after six, 12-18, and 24 hours [70].
In summary, cybersickness symptoms typically last 1-2 hours but can persist for days in extreme cases [24, 54, 75, 81]. Additionally, numerous studies consistently demonstrate an association between habituation or adaptation to cybersickness and repeated exposure to VR applications [24, 54, 75, 81]. There is much less knowledge about the precise models governing the decay over time.
@@ 2.3 Cybersickness and Biophysical Responses
Traditionally, subjective measures such as the SSQ and the FMS have been used to assess the severity of cybersickness. However, these self-reported tools have limitations, including relying on users’ subjective perception, which can be influenced by factors such as mood, prior VR experience, and individual sensitivity. Consequently, there has been growing interest in the use of biophysical measures as objective indicators of cybersickness.
Biophysical measures, such as Electroencephalography (EEG), Electrocardiography (ECG), EDA, and HRV, offer the potential to complement or even replace subjective measures. These physiological signals provide direct insights into the autonomic nervous system’s response to VR stimuli, making them valuable for more accurately assessing the onset and severity of cybersickness. Additionally, biophysical signals may enable the development of real-time systems that predict and mitigate cybersickness, offering a more personalized and precise approach to user comfort.
Recent studies have demonstrated that these biophysical measures can effectively capture the physiological responses associated with cybersickness. For instance, changes in HR, skin conductivity, and stomach activity have been shown to correlate with the severity of cybersickness symptoms [10, 16]. Similarly, EEG measures, particularly changes in brainwave patterns, have been identified as reliable indicators of cybersickness [40, 41, 33]. These findings suggest that biophysical measures could effectively supplement subjective measures like the SSQ, offering more accurate, continuous, and objective assessments of cyber sickness.
@@
[55] employed a combination of EEG, Electromyography (EMG) from the gastrocnemius muscles, and HR to assess motion sickness, finding that EMG parameters were more significant than EEG for predicting motion sickness. In a similar vein, [42] demonstrated that higher power in the gamma band (32Hz) of EEG, coupled with increased heart rate, was associated with heightened cybersickness severity.
@@
The role of multiple physiological signals in predicting cybersickness has been explored in several studies. [10] utilized Electromyography (EOG), EDA, Electrooculography (EGG), ECG, and respiration to study cybersickness. They found that changes in stomach activity, blinking behavior, and breathing patterns could predict cybersickness severity. Likewise, [16] collected data on ECG, EOG, respiratory signals, and skin conductivity from 66 participants and showed that these physiological signals were valuable indicators of cybersickness.
@@
[25] compared HR and EDA signals between high and low cybersickness groups, finding significant differences in HR but no significant differences in breathing patterns or HRV. These results suggest that HR is a more reliable indicator of cybersickness than other measures like respiration rate or HRV.
@@
[40] proposed a method for recognizing VR motion sickness using EEG rhythm energy ratios, achieving high recognition accuracy for individual subjects using a wavelet packet transform. Their method, which combines EEG features with machine learning models such as support vector machines, highlights the growing trend of using computational models to improve the accuracy of motion sickness detection.
@@
[45] explored how HRV and EDA metrics during VR tasks could be used as potential indicators of cybersickness. Similarly, [11] employed a multi-modal approach, incorporating EEG, ECG, EGG, and postural sway to predict motion sickness, finding that EEG-derived features were the most effective in cybersickness prediction, with an accuracy of 93.8%.
@@
More recently, [26] introduced a multimodal deep fusion network that combined HR, EDA, eye movements, and head-tracking data to predict the onset of cybersickness in real-time. These advances reflect a broader trend in the field where most research focuses on ensemble methods and deep learning models to predict cybersickness onset, rather than evaluating and understanding individual contributions of biophysical signals. Similarly, [51] proposed machine learning models using physiological and behavioral features, such as heart rate and body movements, achieving promising results in predicting cybersickness onset.
@@
[34] simulated aircraft motion in VR, and physiological responses were measured, including heart rate, blood pressure, body temperature, and cortisol levels. Their results showed significant increases in heart rate and cortisol levels associated with cybersickness, while blood pressure decreased. The study concludes that HR and cortisol are promising physiological markers for cybersickness assessment and calls for future research on factors specifically affecting disorientation.
In conclusion, while there is substantial research on the relationship between biophysical responses and cybersickness, most studies concentrate on using machine learning techniques, particularly ensemble models, to predict the onset of cybersickness. The integration of multiple biosignals offers promising results, but the individual contributions of each signal remain underexplored, and further research is needed to investigate specifically post-exposure physiological responses.

### v2

@@ Abstract
Virtual Reality (VR) and Augmented Reality (AR) technologies are increasingly being integrated into cognitively demanding work environments, such as aviation, surgery, and complex manufacturing processes. These Extended Reality (XR) technologies offer immersive training opportunities, real-time data visualization, and enhanced decision-making support. However, despite its advantages, prolonged exposure to XR can introduce physiological and cognitive challenges, including motion sickness, fatigue, and stress responses, which may impact performance and well-being [Parsons2018VRChallenges, Rizzo2017clinical]. Understanding how these environments affect users is crucial for optimizing their application in high-stakes professional settings.
@@
A key issue in immersive environments is cybersickness, a form of motion sickness induced by sensory discrepancies between visual and vestibular signals [Reason1975motion]. It manifests as nausea, dizziness, headache, and other autonomic symptoms frequently experienced by VR users [Dennison2016, Keshavarz2015]. Motion sickness in VR can occur under various conditions, including prolonged exposure, high visual flow, latency between head movements and visual updates, and discrepancies between acceleration cues and expected motion [Stanney1997]. The severity and duration of symptoms vary among individuals, with some experiencing mild discomfort that subsides within minutes, while others report persistent symptoms lasting hours after VR exposure [LaViola2000]. Factors such as individual susceptibility, prior VR experience, and display technology play significant roles in determining the intensity of motion sickness effects.
@@
The washout phase of sickness after XR exposure has been studied, though not in great depth. However, it remains a critical aspect to explore further, particularly given its implications for both the robustness of scientific methods [zielasko2024carryOver] and safety in high-stakes environments. For instance, understanding how long symptoms persist and how quickly they subside is essential when considering the use of XR technologies in professions where precise performance is required, such as surgery or aviation. Prolonged exposure to immersive technologies may lead to lingering symptoms that could impair a user’s ability to operate potentially dangerous machinery or perform high-risk tasks. As XR technologies become more prevalent in work environments, studying recovery from sickness is vital to ensure safe and effective usage in critical settings.
@@
Previous studies have demonstrated that cybersickness can elicit a stress response, reflected in autonomic and endocrine changes. Research has shown that individuals experiencing higher discomfort in VR environments exhibit increased cortisol levels, altered Electrodermal Activity (EDA), and Heart Rate (HR) patterns [Bos2005motion, Munafo2017TheVR]. A recent study by [Kim2022] further supports this by demonstrating that cybersickness caused by Head-Mounted Display (HMD)-based VR environments significantly influences physiological responses, including elevated Heart Rate Variability (HRV), altered autonomic nervous system activity, and increased cortisol levels. The physiological stress response is characterized by the activation of the Hypothalamic-Pituitary-Adrenal (HPA) axis and the sympathetic nervous system [McEwen2007]. While the HPA axis regulates cortisol release, the sympathetic nervous system governs rapid autonomic adaptations, manifesting in increased EDA and changes in HR [ulrich2009neural].
@@
Stress has been shown to impact cognitive functions, particularly working memory. While acute stress can enhance cognitive performance under specific conditions [Sandi2013], extensive research suggests that high stress levels impair working memory performance [Schoofs2008, Shields2016]. This impairment may be due to increased cortisol and norepinephrine secretion, which modulate neural networks in the prefrontal cortex, thereby affecting working memory function [Arnsten2009StressPrefrontal].
@@
This study extends previous research by incorporating multiple physiological markers, allowing for a more comprehensive understanding of the motion sickness stress response in VR environments. By examining these effects in a controlled experimental setting, this research contributes to a deeper understanding of the interaction between VR, motion sickness, stress, and cognitive performance. Given the limited and heterogeneous evidence on prolonged post-exposure effects of cybersickness across multiple physiological systems, the present study adopts an exploratory approach to characterize the temporal dynamics of subjective, physiological, and cognitive responses.
@@ 2.1 Temporal Progression of Cybersickness During Exposure
Several studies indicate that cybersickness tends to increase progressively with prolonged exposure. [Kennedy2000duration] observed a linear rise in symptoms among fighter pilots exposed to VR conditions over three hours. Similarly, [Serge2015simulator] reported an upward trend in symptoms over a 16-minute duration, while [Risi2019effects] found a comparable linear increase within the first ten minutes. [McHugh2019] noted a similar trajectory in symptoms over approximately six minutes. [zielasko2018dynamic], who conducted self-reported assessments across two 15-minute conditions, also found indications of a continuous increase.
Longer-duration studies reinforce this trend. [Stanney2003] examined exposure durations of 15, 30, 45, and 60 minutes in a between-subjects design, identifying a sustained increase in cybersickness beyond the 15-minute mark. [Lampton1994] further corroborated this finding, showing that longer exposures correlated with more intense symptoms.
A meta-analysis by [Saredakis2020factors] categorized studies into exposure groups of 0-10 minutes, 10-20 minutes, and beyond 20 minutes. Their findings indicated stable oculomotor symptoms across these groups, whereas nausea and disorientation peaked between 10 and 20 minutes. However, this contrasts with other research suggesting a sustained increase in symptoms beyond 20 minutes.
Some studies suggest adaptation effects over prolonged exposure. In a two-hour experiment, symptoms plateaued after 75 minutes, hinting at potential habituation [Cobb1999]. [Melo2018] found no significant Simulator Sickness Questionnaire (SSQ) [Kennedy1993] score differences between early and later exposure stages, reinforcing the notion of early adaptation. Similarly, studies using the Fast Motion Sickness Questionnaire (FMS) [Keshavarz2011validating] values [McHugh2019, Thorp2022] primarily within the first ten minutes reported adaptation trends.
[Sinitski2018] investigated cybersickness in treadmill-based virtual experiences, noting a significant increase in SSQ disorientation scores up to 15 minutes, followed by an adaptation phase at 45 minutes. Interestingly, nausea peaked at 45 minutes, aligning with [Hakkinen2018time]’s findings, whereas oculomotor symptoms plateaued after 15 minutes. Their data suggest a more complex progression where adaptation occurs alongside symptom escalation.
Other research challenges the notion of a purely gradual increase in symptoms, instead proposing an episodic pattern [Wang2019, Zielasko2021sickness, Teixeira2022unexpected]. Such fluctuations may be difficult to detect due to limitations in sampling rates, which are often constrained by practical feasibility.
@@ 2.2 Temporal Progression of Cybersickness Post Stimulation
Research is sparse on how cybersickness reacts post-stimulation and shows significant interindividual differences in this aspect. For instance, while most subjects experience minimal symptoms after a 10-60 minute break, some continue to suffer from severe symptoms [Singer1998].
A key issue in cybersickness research is the lack of uniformity in post-exposure studies. The need for standard methodologies in this area has been discussed, with suggestions for establishing specific standards [Duzmanska2018].
In a study measuring SSQ 30 minutes after VR exposure, no significant differences were found compared to a pre-study baseline, suggesting rapid recovery in most subjects. However, the disorientation subscale remained significantly higher than pre-experiment levels [Singer1998]. Contrastingly, [Gavgani2017]’s study indicates more prolonged effects of cybersickness. They measured cybersickness before, immediately after, and one, two, and three hours post-exposure over three days. SSQ values were significantly higher post-exposure each day, challenging the notion of rapid recovery suggested by Singer et al. Yet, both studies found significantly lower SSQ values in the first follow-up compared to immediately post-exposure. Further, recovery patterns vary with the VR experiment’s format. In experiments with high SSQ scores, some participants needed over 30 minutes to recover. In contrast, others reported no discomfort just five minutes post-experiment [Tanaka2004virtual]. This observation was mirrored by [Szpak2020exergaming], where most of the participants no longer felt any symptoms 40 minutes after exposure, whereas 14% still felt strong after-effects.
Few studies have examined participants’ well-being the day after the experiment. For example, a study with pilots post-simulator training revealed that 4.6% reported symptoms 24 hours after training [Ungs1987]. Another study with pilots showed that those with intense symptoms had significantly higher values two hours post-experiment compared to a less affected group, but no differences were observed after six, 12-18, and 24 hours [Stoffregen2000].
In summary, cybersickness symptoms typically last 1-2 hours but can persist for days in extreme cases [Ilyas2012, Rebenitsch2016, Tyrrell2018, Zielasko2021sickness]. Additionally, numerous studies consistently demonstrate an association between habituation or adaptation to cybersickness and repeated exposure to VR applications [Ilyas2012, Rebenitsch2016, Tyrrell2018, Zielasko2021sickness]. There is much less knowledge about the precise models governing the decay over time.
@@ 2.3 Cybersickness and Biophysical Responses
Traditionally, subjective measures such as the SSQ and the FMS have been used to assess the severity of cybersickness. However, these self-reported tools have limitations, including relying on users’ subjective perception, which can be influenced by factors such as mood, prior VR experience, and individual sensitivity. Consequently, there has been growing interest in the use of biophysical measures as complementary objective indicators of cybersickness.
Biophysical measures, such as Electroencephalography (EEG), Electrocardiography (ECG), EDA, and HRV, have therefore been explored as potential complementary sources of information alongside subjective measures. These physiological signals provide insights into the autonomic nervous system’s response to VR stimuli, which may reflect physiological changes associated with cybersickness. However, current evidence suggests that no single biophysical measure can yet serve as a reliable standalone indicator of cybersickness across contexts. Additionally, biophysical signals may enable the development of real-time systems that predict and mitigate cybersickness, although such approaches typically rely on multimodal data and context-specific models rather than individual signals alone [Keshavarz2022].
Recent studies have demonstrated that these biophysical measures can capture physiological responses that are associated with cybersickness under specific experimental conditions. For instance, changes in HR, skin conductance level, and gastric activity have been reported to correlate with cybersickness severity in certain setups [Dennison2016, Garcia2019development]. Similarly, several studies have reported associations between specific EEG-derived features (e.g., spectral power in selected frequency bands) and cybersickness-related outcomes [Li2020, Liao2020, Kim2019]. Overall, these findings suggest that biophysical measures may supplement subjective measures such as the SSQ by providing continuous, objective data streams, although their effectiveness appears to be highly dependent on the chosen features, indices, and experimental context. Importantly, reported associations between biophysical measures and cybersickness typically refer to specific indices or features derived from physiological signals (e.g., spectral EEG power in selected frequency bands, heart rate metrics, or statistical properties of the tonic skin conductance level), rather than to the raw signals themselves. Accordingly, several studies emphasize the importance of specifying which indices are examined when interpreting such associations.
@@
[Recenti2021toward] employed a combination of EEG, Electromyography (EMG) from the gastrocnemius muscles, and HR to assess motion sickness, finding that specific EMG-derived parameters were more informative than the selected EEG features for predicting motion sickness. In a similar vein, [Lin2018] demonstrated that increased power in the gamma band (32Hz) of EEG, coupled with elevated heart rate, was associated with heightened cybersickness severity.
@@
The role of multiple physiological signals in predicting cybersickness has been explored in several studies. [Dennison2016] utilized EMG, EDA, Electrooculography (EOG), ECG, and respiration to study cybersickness and reported that specific indices of gastric activity, blink behavior, and breathing patterns were associated with cybersickness severity. Likewise, [Garcia2019development] collected data on ECG, EOG, respiratory signals, and skin conductance from 66 participants and showed that selected features derived from these signals were informative for cybersickness assessment within their experimental setup.
@@
[Islam2020] compared HR and EDA signals between high and low cybersickness groups, finding significant group differences in heart rate, while no significant differences were observed for breathing patterns or the examined HRV indices. These results indicate that heart rate-related measures may be more sensitive to cybersickness-related changes than the specific respiration and HRV indices considered in that study.
@@
[Li2020] proposed a method for recognizing VR motion sickness using EEG rhythm energy ratios, achieving high recognition accuracy for individual subjects using a wavelet packet transform. Their method combines selected EEG features with machine learning models such as support vector machines, highlighting a broader trend toward computational approaches for cybersickness detection.
@@
[Magaki2019] explored how specific HRV and EDA metrics during VR tasks could be used as potential indicators of cybersickness. Similarly, [Dennison2019] employed a multi-modal approach incorporating EEG, ECG, Electrogastrography (EGG), and postural sway, reporting that the selected EEG-derived features yielded the highest predictive performance within their model.
@@
More recently, [Islam2022towads] introduced a multimodal deep fusion network combining HR, EDA, eye movements, and head-tracking data to predict the onset of cybersickness in real time. These advances reflect a broader trend in which research increasingly focuses on multimodal and machine learning-based approaches rather than treating individual biophysical signals as standalone indicators. Similarly, [oh2021machine] proposed machine learning models using physiological and behavioral features, such as heart rate and body movements, achieving promising results for cybersickness onset prediction.
@@
[Kim2022] simulated aircraft motion in VR and measured physiological responses including heart rate, blood pressure, body temperature, and cortisol levels. Their results showed increases in heart rate and cortisol levels associated with cybersickness, while blood pressure decreased. The authors suggest that heart rate and cortisol may serve as useful physiological correlates of cybersickness within simulated flight scenarios and call for future research on factors specifically affecting disorientation.
In conclusion, while there is substantial research examining associations between biophysical responses and cybersickness, most studies emphasize multimodal and machine learning-based prediction approaches. Consistent with recent reviews [Keshavarz2022], there is currently limited evidence that any individual biophysical measure or index can reliably and robustly indicate cybersickness across different VR contexts. The integration of multiple biosignals shows promise, but the individual contributions of each signal remain underexplored, and further research is needed to better understand post-exposure physiological responses.